OpenStudy (anonymous):

jim bought x fishing lures at $4 and y fishing lures at $7, what is the total number of dollars jim spent

OpenStudy (anonymous):

11

OpenStudy (anonymous):

is it 11xy or 11(x+y)

OpenStudy (anonymous):

4x+7y

OpenStudy (anonymous):

or x = 4/11 y = 7/11

OpenStudy (anonymous):

total $ is 11 though

OpenStudy (anonymous):

im sorry, I thought it was "x for $4" not "at $4"

OpenStudy (anonymous):

so yes, it is $ = 4x + 7y